# Beyond Memorization: A Rigorous Evaluation Framework for Medical Knowledge Editing

### v2

@@ D.1 Performance Comparison of Various Editing Paradigms
In real-world scenarios, knowledge updates are always triggered by concrete events: in medicine, for instance, the U.S. Food and Drug Administration’s approval of a novel oncology drug follows positive clinical trial results; in politics, a change in the presidency (e.g., from Biden to Trump) is driven by certified election outcomes; in law, the enactment of a new data-privacy statute typically relies on high-profile regulatory incidents. Under SGR-Edit, these domain-specific event narratives may be supplied by subject-matter experts or, alternatively, sourced automatically using Retrieval-Augmented Generation frameworks (Lewis et al., 2020). By inputting these event narratives, the LLM can produce evidence-grounded rationales that facilitate reliable and transparent knowledge edits in practice.
D.2 Lexical Overlap Analysis of SGR-Edit and RE-Edit
Method (Editing Paradigm) LLaMA-8B LLaMA-3B
ROUGE-L BLEU Lexical avg. ROUGE-L BLEU Lexical avg.
AlphaEdit (SGR-Edit) 0.334 0.149 0.241 0.348 0.166 0.257
AlphaEdit (RE-Edit) 0.165 0.031 0.098 0.170 0.034 0.102
LoRA (SGR-Edit) 0.516 0.370 0.443 0.592 0.459 0.526
LoRA (RE-Edit) 0.211 0.064 0.137 0.242 0.093 0.167
MEMIT (SGR-Edit) 0.363 0.180 0.271 0.356 0.188 0.272
MEMIT (RE-Edit) 0.159 0.031 0.095 0.169 0.036 0.102
Table 5: Comparison of ROUGE-L and BLEU for SGR-Edit and RE-Edit on LLaMA-8B and LLaMA-3B.
@@ D.2 Lexical Overlap Analysis of SGR-Edit and RE-Edit
The results in Table 5 are computed over a human-validated MedMCQAedit subset of high-quality reference explanations (RE) and self-generated rationales (SGR). Since SGR generation can include spurious content, we manually verify each rationale to ensure it faithfully supports the target answer before using it for ROUGE-L and BLEU calculations. This subset covers 64 QA pairs for LLaMA-8B and 77 for LLaMA-3B. These long-form explanations provide rich contextual support for question answering, enabling a precise evaluation of lexical overlap between the injected knowledge and the model’s post-edit outputs. The datasets are available at.
@@
Across all three editing methods: AlphaEdit, LoRA, and MEMIT, SGR-Edit outperforms RE-Edit by a substantial margin in both ROUGE-L and BLEU. For LLaMA-8B, the average lexical overlap (see Lexical avg.) for SGR-Edit is 0.241 (AlphaEdit), 0.443 (LoRA), and 0.271 (MEMIT), compared to just 0.098, 0.137, and 0.095 under RE-Edit. BLEU improvements are equally dramatic: AlphaEdit rises from 0.031 to 0.149, LoRA from 0.064 to 0.370, and MEMIT from 0.031 to 0.180. These gains confirm that SGR—Edit allows evidence-grounded and logical knowledge representation to align more closely with the correct rationale for medical decision making and thus serve as a superior knowledge target.
@@
When comparing across editing methods, LoRA consistently achieves the highest lexical overlap among all paradigms and model sizes. On LLaMA-8B, LoRA SGR-Edit reaches ROUGE-L=0.516 and BLEU=0.370, yielding a text-average of 0.443. This outperforms both AlphaEdit (0.241 lexical avg.) and MEMIT (0.271 lexical avg.), indicating that LoRA’s low-rank adaptation effectively internalizes the rich and context-driven rationales generated by the model. The pattern also holds on LLaMA-3B, where LoRA SGR-Edit achieves a lexical average score of 0.526 versus 0.257 (AlphaEdit) and 0.272 (MEMIT).
@@
In summary, these lexical metrics corroborate our finding that SGR-Edit consistently yields the highest lexical scores, validating its ability to convey deeper understanding rather than superficial information memorization. Furthermore, LoRA emerges as the most effective editing mechanism, capitalizing on the enriched content of self-generated rationales to maximize knowledge integration.
D.3 Qwen2.5–7B Single-Edit Performance
In addition to LLaMA variants, we evaluate Qwen2.5–7B  [^6] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct under the same medical editing protocol. Table 6 reports efficacy, generalization, and retention on MedExQAedit and MedMCQAedit unding single editing.
Consistent with our findings on LLaMA, LoRA and AlphaEdit remain the most effective editing methods for Qwen2.5–7B. LoRA achieves the highest average score on MedMCQAedit (64.2%), driven by strong efficacy (57.4%) and generalization (50.5%), while AlphaEdit closely follows with an average of 60.0%, showing robust efficacy (62.6%) despite slightly lower retention (74.7%). By contrast, GRACE exhibits very low generalization (1.9% on MedExQAedit, 3.2% onMedMCQAedit) despite near-perfect retention (>96%), indicating that this codebook-based and parameter-conserving updates tend to memorize new facts without enabling flexible application to unseen scenarios.
Other parameter-modifying methods, such as ROME and MEMIT, display moderate efficacy (23.2–55.2%) but generalization remains in the 30–43% range, well below the 90%+ figures often reported on simplified, general-domain benchmarks. Similarly, AnyEdit yields balanced retention (>78%) but low generalization (15.8–30.0%) and efficacy (22.6–40.0%). These results confirm the large performance gap noted in RQ1-F1 and RQ1-F2: even state-of-the-art editing methods struggle to exceed 65% combined performance in realistic medical QA settings, exposing critical limitations of existing methods when confronted with domain-specific complexity.
Method Metric MedExQA_edit MedMCQA_edit
LoRA Eff. 61.5 57.4
Gen. 46.2 50.5
Ret. 76.9 84.7
avg. 61.5 64.2
ROME Eff. 36.0 55.2
Gen. 38.0 43.1
Ret. 46.0 82.3
avg. 40.0 60.2
MEMIT Eff. 30.0 54.1
Gen. 34.0 39.2
Ret. 80.0 75.1
avg. 48.0 56.2
GRACE Eff. 36.5 23.2
Gen. 1.90 3.20
Ret. 96.2 97.4
avg. 44.9 41.2
AnyEdit Eff. 40.0 22.6
Gen. 30.0 15.8
Ret. 78.0 85.3
avg. 49.3 41.2
AlphaEdit Eff. 58.0 62.6
Gen. 40.0 42.6
Ret. 78.0 74.7
avg. 58.7 60.0
Table 6: Results on Qwen2.5-7B (Accuracy %). For avg. scores per column: bold is the best, underline is the second best.
D.4 Post-Edit Inference: Two-Step Rationale & Answer vs. One-Step Final Answer
@@
Table 7: Post-Edit accuracy comparison between Final-Answer Only and Rationale + Answer
We compare two prompting strategies on LLaMA-8B over MedMCQAedit: (i) “Rationale + Answer” and (ii) “Final-Answer Only” (see Figure 11). Table 7 presents post-edit accuracy under GTA-Edit, RE-Edit, and SGR-Edit for LoRA, ROME, MEMIT, and AlphaEdit. Overall, providing a chain of thought during post-edit inference yields substantial gains in editing efficacy and generalization, at minimal retention cost.
For LoRA, the average post-edit score climbs from 0.442 to 0.530 (+0.203) when moving to two-step output. Efficacy improves dramatically (GTA-Edit: +0.170, RE-Edit: +0.106, SGR-Edit: +0.049), and Generalization more than twice in the GTA-Edit case (0.220 $\rightarrow$ 0.416). Retention also rises under RE (0.591 $\rightarrow$ 0.727), demonstrating that transparent reasoning reinforces new facts.
ROME shows a mixed pattern: while GTA-Edit average performance drops slightly (0.421 $\rightarrow$ 0.413), RE-Edit and SGR-Edit see gains (+0.053 and +0.084 avg.), reflecting that rationale prompts help when richer contexts are available. MEMIT benefits modestly (+0.006 to +0.067 avg.), with Generalization under SGR especially boosted (0.302 $\rightarrow$ 0.364). AlphaEdit gains across all paradigms (+0.032 to +0.075 avg.), with Retention remaining above 0.789 even after two-step reasoning.
These results confirm that explicit chain-of-thought prompting significantly enhances the model’s ability to apply the injected knowledge, supporting more reliable and interpretable medical knowledge editing.
D.5 Sequential Editing Impact on Common-Domain Capabilities
We split the MMLU benchmark into health-domain and non-health categories: health-domain accuracy is computed over the subjects {anatomy, clinical_knowledge, college_medicine, human_aging, medical_genetics, nutrition, professional_medicine, virology}, while all other subjects are aggregated to compute non-health accuracy.
In addition to the Health vs. Non-Health breakdown (see section 5.4), we further split MMLU into four broad categories—STEM, Humanities, Social Sciences, and Other (Business, Misc.)—to assess general-domain degradation. Table 8 reports category accuracies for LLaMA-8B and LLaMA-3B at 0, 50, and 100 sequential edits.
Model #Edits STEM $\Delta$STEM Humanities $\Delta$Humanities Social Sci. $\Delta$Social Sci. Other $\Delta$Other
@@ D.5 Sequential Editing Impact on Common-Domain Capabilities
Table 8: General-domain category accuracies and absolute drops $\Delta$ on MMLU across sequential medical edits
@@
After 100 sequential medical edits, all common-domain categories exhibit performance declines, with “Other” (which includes business, health and miscellaneous topics) suffering the largest drop (–6.45 pp for LLaMA-8B; –5.86 pp for LLaMA-3B). STEM accuracy decreases by 3.96 pp on 8B and 2.56 pp on 3B, while Social Sciences and Humanities show smaller but nontrivial declines. These results reinforce our RQ4-F2 finding: medical-focused updates not only impair specialized health knowledge but also erode broader general capabilities. The consistent degradation across diverse categories underscores the challenge of maintaining out-of-domain performance during extensive sequential editing.

### v3

@@ D.3 Event‐Driven Rationale Generation for Practical SGR‐Edit
In real‐world scenarios, knowledge updates are always triggered by concrete events: in medicine, for instance, the U.S. Food and Drug Administration’s approval of a novel oncology drug follows positive clinical trial results; in politics, a change in the presidency (e.g., from Biden to Trump) is driven by certified election outcomes; in law, the enactment of a new data‐privacy statute typically relies on high‐profile regulatory incidents. Under SGR-Edit, these domain-specific event narratives may be supplied by subject-matter experts or, alternatively, sourced automatically using Retrieval-Augmented Generation frameworks (Lewis et al., 2020). By inputting these event narratives, the LLM can produce evidence‐grounded rationales that facilitate reliable and transparent knowledge edits in practice.
D.4 External LLM for Rationale Generation
We conducted an evaluation using a subset (100 samples per LLM), comparing performance when rationales are self-generated (SGR-Edit) by the base LLM versus when they are externally generated (EGR-Edit) by DeepSeek-V3. We applied these rationale types to ROME, MEMIT, and AlphaEdit methods, testing on LLaMA-8B and Qwen2.5-7B models.
Method Metric LLaMA-8B Qwen2.5-7B
SGR-Edit EGR-Edit SGR-Edit EGR-Edit
ROME Eff. 0.52 0.49 0.49 0.62
Gen. 0.43 0.49 0.39 0.45
Ret. 0.64 0.75 0.67 0.77
avg. 0.530 0.577 0.517 0.613
MEMIT Eff. 0.52 0.39 0.48 0.56
Gen. 0.39 0.43 0.44 0.43
Ret. 0.64 0.58 0.69 0.73
avg. 0.512 0.467 0.537 0.573
AlphaEdit Eff. 0.52 0.54 0.48 0.54
Gen. 0.28 0.30 0.39 0.41
Ret. 0.82 0.83 0.73 0.68
avg. 0.540 0.557 0.533 0.543
Table 5: Editing performance comparison between SGR-Edit and ECG-Edit
Table 5 shows that EGR-Edit generally yields higher average scores compared to SGR-Edit. Despite this, our primary design philosophy for SGR-Edit was driven by crucial practical considerations, especially for deployment in sensitive domains like medicine. SGR-Edit’s core strength lies in enabling the local base LLM to self-update and refine its knowledge. This approach avoids the inherent security and privacy risks associated with distilling knowledge from or uploading sensitive local medical data to external LLMs/agents in real-world practice. By relying on SGR-Edit, we maintain a self-contained and secure knowledge editing framework, which is paramount in contexts where data privacy and compliance are non-negotiable.
Method Metric Pre-Edit LLaMA-8B LLaMA-3B Qwen-7B
GTA-Edit RE-Edit SGR-Edit GTA-Edit RE-Edit SGR-Edit GTA-Edit RE-Edit SGR-Edit
LoRA Eff. 0 0.466 0.547 0.665 0.144 0.412 0.615 0.440 0.540 0.590
Gen. 0 0.416 0.460 0.503 0.337 0.439 0.492 0.360 0.430 0.440
Ret. 1 0.708 0.727 0.708 0.529 0.658 0.636 0.770 0.730 0.750
avg. 0.530 0.578 0.625 0.337 0.503 0.581 0.523 0.567 0.593
ROME Eff. 0 0.327 0.453 0.528 0.257 0.299 0.439 0.500 0.420 0.590
Gen. 0 0.296 0.403 0.447 0.251 0.406 0.385 0.193 0.510 0.320
Ret. 1 0.616 0.635 0.667 0.561 0.647 0.610 0.560 0.680 0.750
avg. 0.413 0.497 0.547 0.357 0.451 0.478 0.418 0.537 0.553
MEMIT Eff. 0 0.283 0.435 0.520 0.250 0.345 0.455 0.470 0.480 0.575
Gen. 0 0.252 0.380 0.364 0.230 0.315 0.401 0.300 0.400 0.360
Ret. 1 0.648 0.675 0.671 0.485 0.625 0.663 0.570 0.660 0.720
avg. 0.394 0.497 0.518 0.322 0.428 0.506 0.447 0.513 0.552
GRACE Eff. 0 0.366 0.410 0.435 0.230 0.246 0.316 0.230 0.230 0.230
Gen. 0 0.255 0.248 0.317 0.230 0.267 0.316 0.320 0.320 0.320
Ret. 1 0.789 0.758 0.764 0.781 0.747 0.754 0.690 0.690 0.690
avg. 0.470 0.472 0.505 0.414 0.420 0.462 0.413 0.413 0.413
AnyEdit Eff. 0 0.366 0.410 0.435 0.230 0.246 0.316 0.260 0.380 0.370
Gen. 0 0.255 0.248 0.317 0.230 0.267 0.316 0.270 0.160 0.240
Ret. 1 0.789 0.758 0.764 0.781 0.747 0.754 0.830 0.880 0.840
avg. 0.470 0.472 0.505 0.414 0.420 0.462 0.453 0.473 0.483
AlphaEdit Eff. 0 0.439 0.547 0.584 0.326 0.348 0.374 0.510 0.520 0.595
Gen. 0 0.312 0.335 0.366 0.278 0.332 0.348 0.420 0.360 0.362
Ret. 1 0.867 0.795 0.789 0.775 0.743 0.754 0.640 0.700 0.721
avg. 0.539 0.559 0.580 0.460 0.474 0.492 0.523 0.527 0.559
Table 6: Complete results of GTA-Edit, RE-Edit, and SGR-Edit across LLaMA-8B, LLaMA-3B, and Qwen-7B (Accuracy). SGR-Edit achieves the highest average score in nearly all comparisons.
D.5 Performance Comparison of Various Editing Paradigms
As shown in Table 6, relative to GTA-Edit, RE-Edit consistently raises average editing scores on LLaMA-8B by 0.2–8.4 percentage points across methods: ROME sees the largest gain (+8.4 pp), followed by MEMIT (+5.3 pp), LoRA (+4.8 pp), AlphaEdit (+2.0 pp), and AnyEdit (+0.2 pp). These improvements are driven primarily by jumps in Efficacy (up to +12.6 pp for ROME) and Generalization (up to +10.7 pp for ROME), while Retention remains stable above 61% for all methods. Incorporating SGR-Edit yields further average gains of 2.1–7.1 pp over RE-Edit, with MEMIT (+7.1 pp) and LoRA (+4.7 pp) benefiting most. Specifically, LoRA’s combined score climbs from 57.8% to 62.5%, and MEMIT from 44.7% to 51.8%. Even AlphaEdit, which already excels under RE-Edit, improves from 55.9% to 58.0% (+2.1 pp). Critically, these gains in Efficacy and Generalization come with only minor retention trade-offs (e.g., AnyEdit drops from 78.9% to 76.4%), confirming that richer and context-driven rationales enable deeper medical knowledge integration without undue forgetting.
The observed trend of improvement from RE-Edit to SGR-Edit extends robustly to the Qwen-7B model. Across all methods on Qwen-7B, SGR-Edit achieves the highest average editing scores, demonstrating its superiority regardless of the base LLM architecture. Notably, ROME and MEMIT exhibit substantial gains, increasing their average scores by +13.5 pp and +10.5 pp from GTA-Edit to SGR-Edit, respectively, confirming that the enriched SGR target benefits knowledge integration in Qwen-7B as well.
Additionally, we note that the GRACE method exhibited an anomalous pattern across multiple trials: its performance remained nearly identical across all three editing paradigms (i.e., GTA-Edit, RE-Edit, and SGR-Edit). We speculate this is due to GRACE’s reliance on a codebook’s deferral radius, limiting edits to inputs near existing keys, which makes it less sensitive to the content variation. We leave this phenomenon for future investigation.
Figure 13: Interpretability comparison for SGR-Edit and RE-Edit using AlphaEdit, LoRA, and MEMIT. The metric QoR (Quality of Reasoning) represents a normalized score (0 to 1) obtained from an LLM-as-a-Judge (DeepSeek-V3) evaluation, assessing answer accuracy, medical factual accuracy, logical flow, relevance, and completeness.
D.6 Interpretability Analysis of SGR-Edit and RE-Edit
@@
The results in Figure 13 are computed over a human‐validated MedMCQAedit subset of high‐quality reference explanations (RE) and self‐generated rationales (SGR). Since SGR generation can include spurious content, we manually verify each rationale (see Table 4) to ensure it faithfully supports the target answer before using it for ROUGE-L and BLEU calculations. These long‐form explanations provide rich contextual support for question answering, enabling a precise evaluation of lexical overlap between the injected knowledge and the model’s post‐edit outputs.
@@
Across all three editing methods: AlphaEdit, LoRA, and MEMIT, SGR-Edit outperforms RE‐Edit by a substantial margin in both ROUGE‐L and BLEU. For LLaMA‐8B, the average lexical overlap (see Lexical avg.) for SGR‐Edit is 0.241 (AlphaEdit), 0.443 (LoRA), and 0.271 (MEMIT), compared to just 0.098, 0.137, and 0.095 under RE‐Edit. BLEU improvements are equally dramatic: AlphaEdit rises from 0.031 to 0.149, LoRA from 0.064 to 0.370, and MEMIT from 0.031 to 0.180. These gains confirm that SGR—Edit allows evidence‐grounded and logical knowledge representation to align more closely with the correct rationale for medical decision making and thus serve as a superior knowledge target.
@@
When comparing across editing methods, LoRA consistently achieves the highest lexical overlap among all paradigms and model sizes. On LLaMA‐8B, LoRA SGR‐Edit reaches ROUGE‐L=0.516 and BLEU=0.370, yielding a text‐average of 0.443. This outperforms both AlphaEdit (0.241 lexical avg.) and MEMIT (0.271 lexical avg.), indicating that LoRA’s low‐rank adaptation effectively internalizes the rich and context‐driven rationales generated by the model. The pattern also holds on LLaMA‐3B, where LoRA SGR‐Edit achieves a lexical average score of 0.526 versus 0.257 (AlphaEdit) and 0.272 (MEMIT).
We further analyze the overall quality of the generated rationales using the Quality of Reasoning (QoR) Score. This is achieved by prompting the DeepSeek-V3 to evaluate output quality across several dimensions: Answer Accuracy, Medical Factual Accuracy, Logical Flow, Relevance, and Completeness (the prompt are provided in Figure 14). SGR-Edit generally yields a higher Quality Score than RE-Edit, suggesting that using the model’s self-generated context results in more cohesive and contextually rich explanations.
Figure 14: LLM-as-a-judge prompt for predicted rationale (Interpretability) evaluation
@@
In summary, these metrics corroborate our finding that SGR‐Edit consistently yields the highest interpretability, validating its ability to convey a deeper understanding rather than superficial information memorization. Furthermore, LoRA emerges as the most effective editing mechanism, capitalizing on the enriched content of self‐generated rationales to maximize knowledge integration.
D.7 Post-Edit Inference: Two-Step Rationale & Answer vs. One-Step Final Answer
@@
Table 7: Post-Edit accuracy comparison between Final‐Answer Only and Rationale + Answer
We compare two prompting strategies on LLaMA-8B over MedMCQAedit: (i) “Rationale + Answer” and (ii) “Final‐Answer Only” (see Figure 15). Table 7 presents post‐edit accuracy under GTA-Edit, RE-Edit, and SGR-Edit for LoRA, ROME, MEMIT, and AlphaEdit. Overall, providing a chain of thought during post-edit inference yields substantial gains in editing efficacy and generalization, at minimal retention cost.
For LoRA, the average post‐edit score climbs from 0.442 to 0.530 (+0.203) when moving to two‐step output. Efficacy improves dramatically (GTA-Edit: +0.170, RE-Edit: +0.106, SGR-Edit: +0.049), and Generalization more than twice in the GTA-Edit case (0.220 $\rightarrow$ 0.416). Retention also rises under RE (0.591 $\rightarrow$ 0.727), demonstrating that transparent reasoning reinforces new facts.
ROME shows a mixed pattern: while GTA-Edit average performance drops slightly (0.421 $\rightarrow$ 0.413), RE-Edit and SGR-Edit see gains (+0.053 and +0.084 avg.), reflecting that rationale prompts help when richer contexts are available. MEMIT benefits modestly (+0.006 to +0.067 avg.), with Generalization under SGR especially boosted (0.302 $\rightarrow$ 0.364). AlphaEdit gains across all paradigms (+0.032 to +0.075 avg.), with Retention remaining above 0.789 even after two‐step reasoning.
These results confirm that explicit chain‐of‐thought prompting significantly enhances the model’s ability to apply the injected knowledge, supporting more reliable and interpretable medical knowledge editing.
D.8 Sequential Editing Impact on Common‐Domain Capabilities
We split the MMLU benchmark into medical and non‐medical categories: the medical sets are computed over the subjects {anatomy, clinical_knowledge, college_medicine, human_aging, medical_genetics, nutrition, professional_medicine, virology}, which are typically referred to as health-related in the MMLU community. For consistency with our medical evaluation setting, we collectively denote them as the medical subset, while all remaining subjects are aggregated as the non-medical subset.
In addition to the medical vs. non‐medical breakdown (see section 5.4), we further split MMLU into four broad categories—STEM, Humanities, Social Sciences, and Other (Business, Misc.)—to assess general‐domain degradation. Table 8 reports category accuracies for LLaMA‐8B and LLaMA‐3B at 0, 50, and 100 sequential edits.
[…]
@@ D.8 Sequential Editing Impact on Common‐Domain Capabilities
Table 8: General‐domain category accuracies and absolute drops $\Delta$ on MMLU across sequential medical edits
@@
After 100 sequential medical edits, all common‐domain categories exhibit performance declines, with “Other” (which includes business, health, and miscellaneous topics) suffering the largest drop (e.g., –6.45 pp for LLaMA-8B; –5.86 pp for LLaMA-3B; -16.51 for Qwen-7B). STEM accuracy decreases by 3.96 pp on 8B and 2.56 pp on 3B, while Social Sciences and Humanities show smaller but nontrivial declines. These results reinforce our RQ4‐F2 finding: medical‐focused updates not only impair specialized medical knowledge but also erode broader general capabilities. The consistent degradation across diverse categories underscores the challenge of maintaining out‐of‐domain performance during extensive sequential editing.
D.9 Sequential Editing Impact with different Length of Knowledge
To investigate whether the introduction of longer rationales using SGR-Edit would lead to more rapid degradation of general capabilities on MMLU, we conducted experiments using LLaMA-3B as our base LLM. We employed AlphaEdit and MEMIT as sequential editing methods, testing their impact with different rationale token lengths (200-250, 150-200, and 100-150) on MMLU performance after 0, 20, and 50 sequential edits.
[…]
pre-edit 0 60.7 63.7 60.5
AlphaEdit SGR 200-250 50 59.2 -1.5 60.7 -3.0 59.3 -1.2
SGR 150-200 50 58.1 -2.6 60.5 -3.2 58.4 -2.1
SGR 100-150 50 59.6 -1.1 60.7 -3.0 59.8 -0.7
GTA <<10 50 60.0 -0.7 62.3 -1.4 59.9 -0.6
RE 50-100 50 59.3 -1.4 61.0 -2.7 59.4 -1.1
RE 100-50 50 59.0 -1.7 60.4 -3.3 59.0 -1.5
SGR 200-250 20 60.3 -0.4 63.1 -0.6 60.1 -0.4
SGR 150-200 20 60.1 -0.6 62.1 -1.6 60.0 -0.5
SGR 100-150 20 60.6 -0.1 62.8 -0.9 60.4 -0.1
GTA <<10 20 60.4 -0.3 64.1 +0.4 60.8 +0.3
RE 50-100 20 60.6 -0.1 63.9 +0.2 60.3 -0.2
RE 100-50 20 60.6 -0.1 63.4 -0.3 60.3 -0.2
MEMIT SGR 200-250 50 25.5 -35.2 22.8 -40.9 25.8 -34.7
SGR 150-200 50 25.7 -35.0 24.9 -38.8 26.2 -34.3
SGR 100-150 50 25.4 -35.3 22.8 -40.9 25.8 -34.7
GTA <<10 50 25.9 -34.8 26.3 -37.4 26.2 -34.3
RE 50-100 50 24.4 -36.3 26.2 -37.5 24.9 -35.6
RE 100-50 50 26.5 -34.2 25.5 -38.2 24.3 -36.2
SGR 200-250 20 54.8 -5.9 42.3 -21.4 57.0 -3.5
SGR 150-200 20 51.4 -9.3 33.3 -30.4 54.7 -5.8
SGR 100-150 20 54.9 -5.8 45.7 -18.0 57.1 -3.4
GTA <<10 20 48.2 -12.5 33.0 -30.7 50.6 -9.9
RE 50-100 20 43.9 -16.8 27.4 -36.3 48.1 -12.4
RE 100-50 20 54.6 -6.1 40.7 -23.0 56.9 -3.6
Table 9: Performance comparison of AlphaEdit and MEMIT for different knowledge lengths and number of edits on MMLU benchmark.
The experimental results in Table 9 do not indicate that the introduction of longer rationales used for editing leads to a more rapid degradation of general capabilities on MMLU. In contrast, methods like MEMIT exhibit significantly higher degradation regardless of the rationale length, highlighting that the degradation is more dependent on the editing method itself rather than solely on the rationale length.
D.10 Sequential Editing with ROME on External Domains
To further verify the generalization of our findings in RQ4 (Section 5.4), we conducted extended experiments with ROME on LLaMA-3B. As shown in Table 10, similar to MEMIT, the model edited by ROME exhibits severe degradation in both medical and non-medical external domains as the number of edits increases. This contrasts sharply with the relative stability of AlphaEdit (see Figure 6), strongly supporting our conclusion that traditional KE methods fail to conserve external domain knowledge in sequential settings.
Method #Edits Overall Health Non-Health
ROME 0 60.7 63.7 60.5
50 24.1 26.1 23.9
100 22.9 23.9 23.0
Table 10: Sequential editing impact of ROME on MMLU External Domains (LLaMA-3B).
D.11 Sequential Stability Across Methods and Paradigms
In Section 5.5, we observed that MEMIT suffers from model collapse. To investigate whether this issue is method-specific or paradigm-dependent, we extend our analysis to include AlphaEdit, ROME, and MEMIT across all three paradigms (SGR-Edit, RE-Edit, GTA-Edit) on LLaMA-3B.
Table 11 presents the comprehensive results. We observe two key trends:
  • Traditional Methods Collapse: Both ROME and MEMIT exhibit consistent catastrophic collapse (accuracy dropping to $\sim$0.0% on TriviaQA and GSM8K) after 50 edits, regardless of types of target knowledge. This indicates that the instability stems from the parameter-update mechanism of these methods rather than the editing paradigm.
  • AlphaEdit Robustness: In stark contrast, AlphaEdit maintains high stability across all paradigms. Even after 100 edits, SGR-Edit with AlphaEdit retains 62.2% on GSM8K and 27.6% on TriviaQA, significantly outperforming traditional approaches.
Table 11: Sequential editing performance comparison of AlphaEdit, ROME, and MEMIT across different editing paradigms on LLaMA-3B. Red indicates model collapse.
Paradigm #Edits AlphaEdit ROME MEMIT
Hella Trivia Truth GSM Hella Trivia Truth GSM Hella Trivia Truth GSM
SGR-Edit 0 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6
50 70.7 31.6 49.9 64.1 24.7 0.0 51.1 0.0 26.2 0.0 51.5 0.0
100 70.0 27.6 49.9 62.2 23.9 0.0 49.9 0.0 26.5 0.0 48.2 0.0
RE-Edit 0 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6
50 71.0 32.2 48.4 64.1 24.8 0.0 48.6 0.0 26.6 0.0 48.0 0.0
100 70.8 29.3 48.7 62.8 25.7 0.0 48.3 0.0 26.4 0.0 47.9 0.0
GTA-Edit 0 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6 70.4 33.8 49.7 64.6
50 70.9 32.4 49.9 64.3 26.9 0.0 46.3 0.0 30.5 0.0 49.8 0.0
100 70.9 31.2 48.9 63.5 26.2 0.0 48.7 0.0 26.2 0.0 47.2 0.0